(a) Let α : R
3 −→ R
3 be a linear transformation satisfying
α(1, 1, 0) = (1, 2, −1), α(1, 0, −1) = (0, 1, 1) and α(0, −1, 1) = (3, 3, 3).
i. Express (1, 0, 0) as a linear combination of (1, 2, −1), (0, 1, 1) and (3, 3, 3).

ii. Hence find v ∈ R
3
such that α(v) = (1, 0, 0).
(b) Let the map β : R
3 −→ R
3 be defined by
β((a, b, c)) = (a + b + c, −a − c, b)
for any (a, b, c) ∈ R
3
.
i. Show that β is a linear transformation.
ii. Find the kernel of β.
1
Expert's answer
2020-06-25T18:40:05-0400

(a) i.

"(1, 0, 0)=c_1(1, 2, \u22121)+c_2(0, 1, 1)+c_3(3, 3, 3)"

"c_1+3c_3=1""2c_1+c_2+3c_3=0""-c_1+c_2+3c_3=0"

"c_1+3c_3=1""3c_1=0""-c_1+c_2+3c_3=0"

"c_1=0""c_2=-1""c_3={1\over 3}"

"(1, 0, 0)=(0)(1, 2, −1)+(-1)(0, 1, 1)+({1\over 3})(3, 3, 3)"

ii.

"\text{v}=(0)(1, 1, 0)+(-1)(1, 0, -1)+({1\over 3})(0, -1, 1)"

"\text{v}=(-1, -{1\over 3}, {4\over 3})"

(b) i.


"\u03b2((a, b, c)) = (a + b + c, \u2212a \u2212 c, b)"

Let "\text{x}=(a_1, b_1,c_1), \text{y}=(a_2, b_2,c_2), a_1, b_1,c_1,a_2, b_2,c_2\in \R"



"\beta(\text{x}+\text{y})=\beta((a_1, b_1,c_1)+(a_2, b_2,c_2))="


"=\beta((a_1+a_2, b_1+b_2,c_1+c_2)="


"=(a_1+a_2+b_1+b_2+c_1+c_2, -(a_1+a_2)-(c_1+c_2), b_1+b_2)="

"=(a_1+b_1+c_1, -a_1-c_1,b_1)+(a_2+b_2+c_2,-a_2-c_2,b_2)="

"=\beta((a_1, b_1, c_1))+\beta((a_2, b_2, c_2))=\beta(\text{x})+\beta(\text{y})"

Let "r\in \R"


"\beta(r\text{x})=\beta(r(a_1, b_1,c_1))=\beta((ra_1, rb_1,rc_1))="

"=(ra_1+rb_1+rc_1, -ra_1-rc_1,rb_1)="

"=r(a_1+b_1+c_1, -a_1-c_1,b_1)=r\beta(\text{x})"

Therefore "\beta" is a linear transformation.  


ii.

System of linear equations associated to the implicit equations of the kernel, resulting from equalling to zero the components of the linear transformation formula.


"a+b+c=0""-a-c=0""b=0"

The system has infinitely many solutions:


"a=-c""b=0""c=arbitrary"


"A=\begin{pmatrix}
 1 & 1 & 1 \\
 -1 & 0 & -1 \\
 0 & 1 & 0
\end{pmatrix}"

"\begin{pmatrix}
 1 & 1 & 1 \\
 -1 & 0 & -1 \\
 0 & 1 & 0
\end{pmatrix}\xmapsto{R_2=R_2+R_1}\begin{pmatrix}
 1 & 1 & 1 \\
 0 & 1 & 0 \\
 0 & 1 & 0
\end{pmatrix}"

"\begin{pmatrix}
 1 & 1 & 1 \\
 0 & 1 & 0 \\
 0 & 1 & 0
\end{pmatrix}\xmapsto{R_3=R_3-R_2}\begin{pmatrix}
 1 & 1 & 1 \\
 0 & 1 & 0 \\
 0 & 0 & 0
\end{pmatrix}"

"\begin{pmatrix}
 1 & 1 & 1 \\
 0 & 1 & 0 \\
 0 & 0 & 0
\end{pmatrix}\xmapsto{R_1=R_1-R_2}\begin{pmatrix}
 1 & 0 & 1 \\
 0 & 1 & 0 \\
 0 & 0 & 0
\end{pmatrix}"

The solution can be written in the vector form:


"\begin{pmatrix}
 a \\
 b \\
c
\end{pmatrix}=\begin{pmatrix}
 -1 \\
 0 \\
1
\end{pmatrix}c"

Therefore the kernel of "\beta" has a basis formed by the set


"\bigg\{\begin{pmatrix}
 -1 \\
 0 \\
1
\end{pmatrix}\bigg\}"